line with a slope of -3 passing through (4,-5)

Found 2 solutions by ankor@dixie-net.com, Theo:
Answer by ankor@dixie-net.com(22740)

line with a slope of -3 passing through (4,-5)
Using the point/slope formula y - y1 = M(x - x10
Substituting
y - (-5) = -3(x - 4)
y + 5 = -3x + 12
y = -3x + 12 -5
y = -3x + 7 is the line passing thru 4, -5)

Answer by Theo(13342)

the line has a slope of -3 passing through (4,-5)

slope intercept form of a straight line is y = mx + b.
m is the slope]b is the y-intercept.

if the slope is -3, the general form becomes y = -3 * x + b

to solve for b, you replace y with -5 and x with 4 to get -5 = -3 * 4 + b

simplify to get -5 = -12 + b
add 12 to both sides of the equation to get 7 = b

the equation becomes y = -3 * x + 7.

here's what it looks like on a graph.



the y-intercept is 7.
that's the value of y when x is 0.
the line passes through the point (4,-5).
the slope is -3.